OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x+y=7 9x+___y=___

OpenStudy (anonymous):

@Directrix plz help

OpenStudy (anonymous):

i think it is 9x+10y=16

Directrix (directrix):

Click on the attachment to see how coincident lines look.

OpenStudy (anonymous):

wait no give me a sec plz

Directrix (directrix):

They are the same line so the equations will be the same although appear different because the same number has been multiplied by every term of the other equation.

OpenStudy (anonymous):

so 2x3 and 2x7?

OpenStudy (anonymous):

or 3x3 and 3x7

Directrix (directrix):

3x +y = 7 9x +___y=___ ============ What do you multiply 3x by to get 9x?

OpenStudy (anonymous):

i would multiply 3x3 to get 9

OpenStudy (anonymous):

so i would have 9x+3y=21

Directrix (directrix):

3x +y = 7 Okay, then multiply each of the 3 terms above by 3 to get the equation of the coincident line.

Directrix (directrix):

>>so i would have 9x+3y=21 Correct

OpenStudy (anonymous):

thanks